What is the magnitude of a point charge chosen so that the electric field intensity E has a magnitude of 3.30 N/C at a distance of 76cm. what do you mean by point charge?


1
Expert's answer
2021-01-29T20:02:52-0500
"E=k\frac{q}{r^2}\\3.3=8.99\cdot10^9\frac{q}{0.76^2}\\q=2.12\cdot10^{-10}\ C"

The point charge is an electric charge considered to exist at a single point, and thus having neither area nor volume.